Find the slope of the line: 2x + 15y = 29
a. m=2/15
b. m= -2/15

Answer by csr212(7)

m= -2/15
2x+15y=29
Since 2x does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 2x from both sides.
15y=-2x+29
Divide each term in the equation by 15.
15y/15 = -(2x/15)+29/15
Simplify the left-hand side of the equation by canceling the common factors.
y=-(2x/15)+29/15
To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept.
y=mx+b
Using the y=mx+b formula,- 2/15 and b= 29/15